Question 2

Cullumber Ltd. purchased a new machine on April 4, 2014, at a cost of $179,200. The company estimated that the machine would have a residual value of $16,000. The machine is expected to be used for 10,200 working hours during its four-year life. Actual machine usage was 1,400 hours in 2014; 2,200 hours in 2015; 2,400 hours in 2016; 2,200 hours in 2017; and 2,000 hours in 2018. Cullumber has a December 31 year end.

Calculate depreciation for the machine under each of the following methods: (Round expense per unit to 2 decimal places, e.g. 2.75 and final answers to 0 decimal places, e.g. 5,275.)

(1) Straight-line for 2014 through to 2018.
2014 expense $enter a dollar amount

2015 expense $enter a dollar amount

2016 expense $enter a dollar amount

2017 expense $enter a dollar amount

2018 expense $enter a dollar amount



(2) Diminishing-balance using double the straight-line rate for 2014 through to 2018.
2014 expense $enter a dollar amount

2015 expense $enter a dollar amount

2016 expense $enter a dollar amount

2017 expense $enter a dollar amount

2018 expense $enter a dollar amount



(3) Units-of-production for 2014 through to 2018.
2014 expense $enter a dollar amount

2015 expense $enter a dollar amount

2016 expense $enter a dollar amount

2017 expense $enter a dollar amount

2018 expense $enter a dollar amount

Which method results in the highest depreciation expense over the life of the asset? Highest net income? Highest cash flow?

select a method that results in the highest depreciation expense over the life of the asset

Straight-line methodDiminishing-balance methodUnits-of-production methodAll three methods are the sameNo Impact



Which method results in the highest net income?

select a method that results in the highest net income

Straight-line methodDiminishing-balance methodUnits-of-production methodAll three methods are the sameNo Impact



Which method results in the highest cash flow?

select a method that results in the highest cash flow

Straight-line methodDiminishing-balance methodUnits-of-production methodAll three methods are the sameNo Impact
Answer #1
Straight-Line Depreciation Schedule
Depreciation for the Year
Asset Depreciable Useful Life Depreciation Accumulated Book
year Cost Cost Expense Depreciation Value
2014 $179,200.00
2014 $163,200.00 / 4*9/12 = $30,600.00 $30,600.00 $148,600.00
2015 $163,200.00 /                     4 = $40,800.00 $71,400.00 $107,800.00
2016 $163,200.00 /                     4 = $40,800.00 $112,200.00 $67,000.00
2017 $163,200.00 /                     4 = $40,800.00 $153,000.00 $26,200.00
2018 $163,200.00 / 4*3/12 = $10,200.00 $163,200.00 $16,000.00
Depreciable cost =179200-16000
Depreciation per Unit = (Original Cost- Salvage Value)/ Estimated Unit
=(179200-16000)/10200 Unit=$16
Units-of-Production Depreciation Schedule
Depreciation for the Year
Asset Depreciation Number of Depreciation Accumulated Book
year Cost Per Unit Unit Expense Depreciation Value
2014 $179,200
2014 16 x 1400 = $22,400 $22,400 $156,800.00
2015 16 x 2200 = $35,200 $57,600 $121,600.00
2016 16 x 2400 = $38,400 $96,000 $83,200.00
2017 16 x 2200 = $35,200 $131,200 $48,000.00
2018 16 x 2000 = $32,000 $163,200 $16,000.00
Dimnishing balance
Depreciation for the Year
Asset Book DDB Depreciation Accumulated Book
Date Cost Value Rate Expense Depreciation Value
2014 $179,200
2014 $179,200 x 50% = $67,200 $67,200 $112,000
2015 $112,000 x 50% = $56,000 $123,200 $56,000
2016 $56,000 x 50% = $28,000 $151,200 $28,000
2017 $28,000 = $12,000 $163,200 $16,000
2018 (28000-16000)
Depreication Rate= 1/ Estimated Life X 2= 1/4X2= 50%